• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Punjabi

    Log entry time 23:48:58 on June10,2000

    Entry number 43407

    keyword=Shift Summary

    Started shift with run 1901

    Beam is bad, triping of every couple of minuites.


    18:20 beam became more stable for some time, but now it is tripping again often.


    19:30 beam has been stable since past 30 minuites.

    20:30 MCC called. Radiation trip. No beam for few minutes. Stop run 1914.

    20:49 Electron Q3 quenched again. Asked for an access

    21:20 Q3 is reset and now ramoing up to 1600 amp to recycle.

    21:30 no beam because of water flow fault to beam diffuser. Mark Stevens (Tech on call) is comming to see what can be done.

    23:50 Mark Stevens and Rusty Salmons are still working on the water pump for beam diffuser. Estimate 1 to 2 hours.

    Start Time.....Run Number....Events.....Current(ua)......Charge(mc)

    16:03 1901 1.3 M 61.5 56
    16:33 1902 1.3 M 60 54.7
    17:00 1903 1.3 M 59.5 54.9
    17:20 1904 1.3 M 59.1 54.8
    17:36 1905 1.3 M 58.9 55.1
    17:54 1906 1.3 M 58.5 55.2
    18:12 1907 1.3 M 58.3 55.3
    18:35 1908 1.3 M 58.0 55.5
    18:54 1909 1.3 M 57.3 55.7
    19:11 1910 1.3 M 57.2 55.7
    19:29 1911 1.3 M 57.0 55.7
    19:48 1912 1.3 M 57.2 55.7
    20:08 1913 1.3 M 57.2 55.8
    20:27 1914 0.2 M 57.2 6.7
    20:34 1915 1.1 M 57.4 54.3